OpenStudy (anonymous):

sin2x/1+cos2x=tanx how do i do this

OpenStudy (anonymous):

sin2x=2sinxcosx

OpenStudy (anonymous):

what you need?

OpenStudy (anonymous):

i dont know how sin2x/1+cos2x equals tanx

OpenStudy (anonymous):

trig identities

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sin^2(x)+cos^2(x)=1

OpenStudy (anonymous):

sin^2(x)= (1-cos(2x))/2

OpenStudy (jhannybean):

maybe it's \[\frac{\sin(2x)}{1+cos(2x)}\]

OpenStudy (anonymous):

sin(2x)=2*sin(x)cos(x) cos^2(x)=1/2(1+cos(2x) cos(2x)=2*cos^2(x)-1

OpenStudy (anonymous):

sin^2(x)+cos^2(x)=1 multply the eq by 2 and cos(x) 2cos(x)sin^2(x)+2cos(x)cos^2(x)=2cos(x) sin(2x)*sin(x)+2cos(x)cos^2(x)=2cos(x) sin(2x)*sin(x)+2cos(x)(1/2+cos(2x)/2)=2cos(x) divide by cos (x) sin(2x)*tan(x)+1+cos(2x)=2 take the 1 to the other side sin(2x)*tan(x)+cos(2x)=1

OpenStudy (anonymous):

look at it might help....it needs some thinking

OpenStudy (anonymous):

...\[\frac{ \sin(2x) }{ 1 }+ \cos(2x)= \tan(x)\] if this the equation like that ..... then something wrong here. i put value for x and the amswer are not the same....